### v3

# Coboundaries of 3-IETs

Abstract.
In this note, we investigate the coboundaries of interval exchange transformations of 3 intervals (3-IETs). More precisely, we show that a differentiable function with absolutely continuous derivative with bounded variation, whose integral and integral of its derivative is 0, is a coboundary for typical 3-IET if and only if the values at the endpoints of the domain are zero. We also show the existence of rare counterexamples for both cases of possible values at the endpoints of the interval. We obtain our result by studying the properties of associated skew products.
1. Introduction
One of the most classical problems in the theory of dynamical systems is investigating whether for a given measure preserving transformation $T$ on a standard probability Borel space $(X,\mu)$, a given 00-mean function $f\in L^{1}(X,\mu)$ is a coboundary, i.e. there exists $g\in L^{1}(X,\mu)$ (or with certain regularity conditions) such that the cohomological equation is satisfied:
$f=g\circ T-g.$
This problem has been studied from many points of view and for many classes of dynamical systems, among them, the interval exchange transformations were studied in [MMY05] and in [BT24], while [For97] studied the analogous problem for flows on compact surfaces. The main application of this paper is solving the cohomological equation to determine whether a skew product $T_{f}$ on $X\times\mathbb{R}$ given by
$T_{f}(x,r):=(Tx,r+f(x))$
is ergodic with respect to the product measure $\mu\otimes Leb_{\mathbb{R}}$. It is a classical fact that if $f$ is a coboundary, then the orbit of the point $(x,r)$ is contained in a shifted graph of $f$. Consequently, $T_{f}$ is not ergodic. If one can prove that $T_{f}$ is ergodic, then $f$ is not a coboundary.
We study this problem for the following setup: we consider $X=[0,1)$ with Lebesgue measure ${\rm Leb}$, a symmetric interval exchange transformation of 3 intervals (see Section 2 for definitions) and $f\in C^{1+AC}([0,1))$, where $C^{1+AC}(\cdot)$ denotes the set of differentiable functions whose derivative is absolutely continuous and of bounded variation. We consider two cases, depending on the values of $f$ at the endpoints of the interval. First, we assume that the function $f$ vanishes at the endpoints of the interval $[0,1)$:
Theorem 1.
Let $f\in C^{1+AC}([0,1))$ such that $f(0)=\lim_{x\to 1}f(x)=0$ with $\int f=0$, then $f$ is a coboundary for almost every symmetric 3-IET $T$. In particular, the skew product $T_{f}$ is not ergodic. However, there exists a dense subset of 3-IETs, for which the skew product $T_{f}$ is ergodic.
On the contrary, we have the following result when $f$ does not vanish at 00 and $1$.
Theorem 2.
Let $f\in C^{1+AC}([0,1))$ with $\int f=0$ such that $f(0)=\lim_{x\to 1}f(x)\neq 0$, then the skew product $T_{f}$ is ergodic for almost every symmetric 3-IET $T$. In particular, $f$ is not a coboundary for $T$.
Similarly to the Theorem 1, here we also have a small family of maps, which exhibit different behavior than the typical case.
Proposition 3.
If $f(0)=\lim_{x\to 1}f(x)\neq 0$, then there exists a dense family of minimal $3$-IETs for which $f$ is a coboundary. In particular, the skew product $T_{f}$ is not ergodic.
[…]
It is worth noting that both sine and cosine are coboundaries for rotations. Indeed, if $\alpha$ is an irrational number and $R(x):=x+\alpha\mod 1$, then there exists a continuous function $g:\mathbb{R}/\mathbb{Z}\to\mathbb{R}$, such that $f(x)=g(Rx)-g(x).$ To be more precise in finding $g$, first observe that by the difference of cosines identity, we can write
$\begin{split}\sin(2\pi x)&=\frac{\cos(2\pi(x+\alpha))-\cos(2\pi(x-\alpha))}{-2\sin(2\pi\alpha)}\\ &=\frac{\cos(2\pi(x+\alpha))+\cos(2\pi x)}{-2\sin(2\pi\alpha)}-\frac{\cos(2\pi x)+\cos(2\pi(x-\alpha))}{-2\sin(2\pi\alpha)}.\end{split}$
Hence $f(x)=\sin(2\pi x)$ is a coboundary with solution to the cohomological equation:
$g(x):=\frac{\cos(2\pi x)+\cos(2\pi(x-\alpha))}{-2\sin(2\pi\alpha)},\ x\in[0,1).$
Moreover, by differentiating both sides, $\cos(2\pi x)$ is a coboundary with solution of the cohomological equation $-\frac{1}{2\pi}g^{\prime}(x)$. This contrasts with Theorem 2.
Our result directly relates to the question asked by Chaika and Robertson in [CR19]. The authors in [CR19, Question 1.5] ask whether for a typical IET in at least three intervals, the skew product is ergodic when considering the cocycle $f(x)=\cos(2\pi x),$. Thus, we provide a partial answer in the affirmative to this question using Theorem 2:
Corollary 4.
For almost every 3-IET $T$ on the interval $[0,1)$ and $f(x)=\cos(2\pi x)$, the skew product $T_{f}$ is ergodic.
In contrast, from Theorem 1 we deduced the following:
Corollary 5.
For almost every 3-IET $T$ on the interval $[0,1)$ and $f(x)=\sin(2\pi x)$, the skew product $T_{f}$ is not ergodic. However, there exists a dense subset of 3-IET for which $T_{f}$ is ergodic.
In Section 2, we recall basic definitions and facts concerning IETs, rotations, and skew products. In Section 3, we state the essential value criterion for the ergodicity of skew products (Theorem 11). Then, we pass to the proofs of the main theorems. In Section 4, we outline the proof of Theorem 1. We used the results of [MMY05] and [Frą00]. Finally, in Section 5, we present the proof of Theorem 2 using the classical Denjoy-Koksma inequality, we prove Proposition 3 at the end of the paper.
Acknowledgments
The authors thank Krzysztof Frączek for suggesting the problem and Jon Chaika for his continuous support. This research was partially supported by the Narodowe Centrum Nauki Grant 2022/45/B/ST1/00179.
2. Rotations and interval exchange transformations
Given $\alpha\in\mathbb{R}$, we denote by $R:[0,1)\to[0,1)$ the map $x\mapsto x+\alpha\mod 1$. This map is called rotation by $\alpha$ on $[0,1)$. It is a special case of a more general class of maps called interval exchanges. For more information about this class of maps, we refer the reader to [Via06, Yoc10].
Let $\mathcal{A}$ be an alphabet of $d$ symbols, and let $\lambda\in\mathbb{R}_{+}^{\mathcal{A}}$ be a vector of positive entries. Given two bijective functions $\pi_{t}:\mathcal{A}\to\{1,\dots,d\}$ and $\pi_{b}:\mathcal{A}\to\{1,\dots,d\}$, let $\pi$ be the permutation of symbols in $\mathcal{A}$ defined by
$\pi=\begin{pmatrix}\pi_{t}^{-1}(1)&\dots&\pi_{t}^{-1}(d)\\ \pi_{b}^{-1}(1)&\dots&\pi_{b}^{-1}(d)\\ \end{pmatrix}.$
Let $|\lambda|=\sum_{a\in\mathcal{A}}\lambda_{a}$, $I=[0,|\lambda|)$ and let $I_{a}\subset I$ be the interval
[$\left[\sum_{\{b:\pi_{t}^{-1}(b)<\pi_{t}^{-1}(a)\}}\lambda_{b},\sum_{\{b:\pi_{t}^{-1}(b)\leq\pi_{t}^{-1}(a)\}}\lambda_{b}\right).$
The vector $\lambda$ and the permutation $\pi$ determine the interval exchange transformation on $d$ intervals $T_{\lambda,\pi}:I\to I$ by
$T_{\lambda,\pi}(x)=x+\sum_{\{b:\pi_{b}^{-1}(b)<\pi^{-1}_{b}(a)\}}\lambda_{b}-\sum_{\{b:\pi_{t}^{-1}(b)<\pi_{t}^{-1}(a)\}}\lambda_{b},\quad\text{if }x\in I_{a}.$
We will write $T$ instead of $T_{\lambda,\pi}$ if $\lambda$ and $\pi$ are given. Also, we say $d$-IET instead of “IET in $d$ intervals”.
The (possible) discontinuities of $T_{\lambda,\pi}$ are of the form
$\xi_{\beta}=\sum_{\pi_{t}(\gamma)<\pi_{t}(\beta)}\lambda_{\gamma}$
and denote the set of discontinuities by $D(T_{\lambda,\pi})=\{\xi_{\beta}:\beta\in\mathcal{A}\}$.
An IET $T=T_{\lambda,\pi}$ is said to have the Keane property if for all $m\in\mathbb{N}$ and for all pairs of symbols $\beta,\gamma\in\mathcal{A}$, then
$T^{m}\xi_{\beta}\neq\xi_{\gamma}.$
This work focuses on 3-IETs where the permutation is of the form
$\pi=\begin{pmatrix}A&B&C\\ C&B&A\end{pmatrix},$ (1)
and $2$-IETs (rotations) with
$\pi=\begin{pmatrix}A&B\\ B&A\end{pmatrix}.$
2.1. Induced maps of 3-IETs are rotations
Let $T=T_{\lambda,\pi}$ be an IET and $f:[0,|\lambda|)\to\mathbb{R}$ be a real function. We denote $T_{f}:[0,|\lambda|)\times\mathbb{R}\to[0,|\lambda|)\times\mathbb{R}$ the skew product map $(x,t)\mapsto(Tx,t+f(x))$.
For every fixed 3-IET, we consider a first return map to a properly chosen subinterval $J\subset[0,|\lambda|)$, where $J$ is either $[0,\lambda_{A}+\lambda_{B})$ or $[0,\lambda_{B}+\lambda_{C})$, see Cases 1 and 2 below. This corresponds to the classical Rauzy-Veech induction step, see ex. [Via06] for more details.
We denote $R_{\widetilde{f}}:J\times\mathbb{R}\to J\times\mathbb{R}$ the first return map of the skew product $T_{f}$ to $J\times\mathbb{R}$. Then $R_{\widetilde{f}}:(x,t)\mapsto(T|_{J}(x),t+\widetilde{f}(x)).$
Let $\lambda=(\lambda_{A},\lambda_{B},\lambda_{C})$ be a vector of positive real entries and let $\pi$ be the permutation in Equation 1.
Case 1: $\lambda_{A}>\lambda_{C}$. Let $T|_{J}$ be the first return map of $T$ to $J=[0,\lambda_{A}+\lambda_{B})$. By careful calculations, we can see that $T|_{J}$ is a rotation by $\alpha=\lambda_{B}+\lambda_{C}$ modulo $|J|=\lambda_{A}+\lambda_{B}$. The new cocycle function for the first return map is defined by
$\widetilde{f}(x)=\left\{\begin{array}[]{cll}f(x)&\text{ if }&0\leq x<\lambda_{A}-\lambda_{C}\\ f(x)+f(x+\lambda_{B}+\lambda_{C})&\text{ if }&\lambda_{A}-\lambda_{C}\leq x<\lambda_{A}\\ f(x)&\text{ if }&\lambda_{A}\leq x<\lambda_{A}+\lambda_{B}\end{array}\right.$ (2)
Case 2: $\lambda_{A}<\lambda_{C}$. Let $J=[0,\lambda_{B}+\lambda_{C})$ be the interval of induction. The first return map of $T$ to $J$ is the rotation by $\alpha=\lambda_{C}-\lambda_{A}$ modulo $|J|=\lambda_{B}+\lambda_{C}$. Similarly, the cocycle function for the first return map is the function
$\widetilde{f}(x)=\left\{\begin{array}[]{cll}f(x)+f(x+\lambda_{B}+\lambda_{C})&\text{ if }&0\leq x<\lambda_{A}\\ f(x)&\text{ if }&\lambda_{A}\leq x<\lambda_{B}+\lambda_{C}\end{array}\right.$ (3)
Remark 6.
If $f(0)=\lim_{x\to|\lambda|}f(x)=0$ and $f$ is absolutely continuous, then: In Case 1 (Equation 2) $\widetilde{f}$ is not continuous, but it is piecewise absolutely continuous. The intervals of absolute continuity are $[0,\lambda_{A}-\lambda_{C})$ and $[\lambda_{A}-\lambda_{C},\lambda_{A}+\lambda_{B})$. These two intervals are exchanged by the rotation $T|_{J}$.
In Case 2 (Equation 3) the cocycle $\widetilde{f}$ is absolutely continuous as well. This absolute continuity can be extended to $\mathbb{T}$.
Remark 7.
When proving that a function is a coboundary, we often demonstrate that the induced cocycle is a coboundary rather than directly working with the original cocycle. This is enough for our purposes. Consider an ergodic measure preserving system $(X,\mathcal{B},\nu,T)$ that is invertible. Let $\widetilde{T}:=T|_{Y}$ be the first return map of $T$ to $Y$ for some $Y\subset X$ of $\nu$-positive measure. Let $\widehat{f}:Y\to\mathbb{R}$ be given by
$\widehat{f}(x):=S_{r(x)}f(x),$
where $r(x)$ is the first return time of $x$ to $Y$ via $T$. If $\widehat{h}:Y\to\mathbb{R}$ is a solution of the cohomological equation
$\widehat{f}(x)=\widehat{h}(\widetilde{T}x)-\widehat{h}(x),$
then for every $x\in Y$ and $j=0,\ldots,r(x)-1$ by putting
$h(T^{j}x):=\begin{cases}\widehat{h}(x)&\text{ for }j=0,\\ h(T^{j-1}(x))+f(T^{j-1}(x))&\text{otherwise},\end{cases}$
we get that $f(y)=h(Ty)-h(y)$ for every $y\in\bigcup_{n=0}^{\infty}T^{n}Y$. By ergodicity of $T$, the set $\bigcup_{n=0}^{\infty}T^{n}Y$ has full measure and the equation $f(y)=h(Ty)-h(y)$ holds $\nu$-a.e.
Note that if $f$ is piecewise absolutely continuous, so is $\widetilde{f}$. Pask in [Pas90, 2.11. Theorem] proved that for every irrational rotation $R$ and piecewise absolutely continuous function $f$ with non-zero total sum of jumps, the skew product $R_{f}$ is ergodic. Hence, we get the following fact because of Remark 7.
Proposition 8.
For every ergodic 3-IET $T$ on $[0,1)$, if $f\in C^{1}([0,1))$ is such that $f(0)\neq\lim_{x\to 1}f(x)$, then the skew product $T_{f}$ is ergodic.
2.2. Continued fractions
[…]
$a_{0}+\frac{1}{a_{1}+\frac{1}{a_{2}+\frac{1}{a_{3}+\frac{1}{\ddots}}}}$
with $a_{0}\in\{0\}\cup\mathbb{N}$, $a_{1},a_{2},a_{3},\dots\in\mathbb{N}$ be its continued fraction expansion. The coefficients $a_{i}$ are called the partial quotients and the fractions of the form
$\frac{p_{n}}{q_{n}}=a_{0}+\frac{1}{a_{1}+\frac{1}{a_{2}+\frac{1}{\ddots\frac{1}{a_{n-1}+\frac{1}{a_{n}}}}}}$
are called the partial fractions of $\alpha$. We denote by $\|x\|$ the distance of $x$ to the nearest integer and $\{x\}$ the fractional part of $x$. We refer the reader to [Khi97]and [EW10, Chapter 3] for more information on the properties of continued fraction expansions.
We say that $\alpha$ an irrational number is of Roth type if for all $\epsilon>0$ the growth rate of the partial quotients is of the form
$a_{n}=\text{O}(q_{n}^{\epsilon}).$
An equivalent definitions is: for all $\epsilon>0,$ there exists a positive constant $C_{\epsilon}$ such that for all $n\in\mathbb{N}$
$q_{n}^{1+\epsilon}\|q_{n}\alpha\|>C_{\epsilon}.$
The set of all Roth type irrational numbers have full measure.
We note that [MMY05] generalized this property to IETs with more than two intervals. We do not need the definition of Roth type $d$-IETs with $g\geq 3$. However, Roth type 2-IETs are rotations by an irrational number of Roth type.
There is a residual set of irrational numbers that are not of Roth type. Let $k\in\mathbb{N}$ be a natural number and denote $S_{k}^{0}$ the set of irrational numbers $\alpha$ such that the sequence of denominators of the continued fractions $(q_{n})_{n\in\mathbb{N}}$ satisfy that
$\liminf_{n\to\infty}q_{n}^{k+1}\|q_{n}\alpha\|=0.$
2.3. Rokhlin towers for rotations
With the properties of continued fractions defined in Section 2.2, we can construct a family of Rokhlin towers for the interval $[0,1)$ and the rotation by $\alpha$. The Rokhlin tower is built out of two towers, one small of width $\|q_{n+1}\alpha\|$ and height $q_{n}$, and the other tower is wider and very tall of width $\|q_{n}\alpha\|$ and height $q_{n+1}$. For each $n\in\mathbb{N}$ we have that:
If $n$ is even the basis of the Rokhlin tower are the intervals $I_{n}=[-\|q_{n}\alpha\|,0)$ and $\tilde{I}_{n}=[0,\|q_{n+1}\alpha\|)$. The levels of the smaller tower are intervals of the form $R^{j}\tilde{I}_{n}$ with $j\in\{0,\dots,q_{n}-1\}$. The levels of the larger tower are intervals of the form $R^{k}I_{n}$ with $k\in\{0,\dots,q_{n+1}-1\}$.
If $n$ is odd the basis are $\tilde{I}_{n}=[-\|q_{n+1}\alpha\|,0)$ and $I_{n}=[0,\|q_{n}\alpha\|)$. The levels of the tower are intervals of the form $R^{j}\tilde{I}_{n}$ with $j\in\{0,\dots,q_{n}-1\}$ and $R^{k}I_{n}$ with $k\in\{0,\dots,q_{n+1}-1\}$.
[…]
(a) Case $n$ is even
[…]
(b) Case $n$ is odd
Figure 1. Rokhlin towers for a rotation by $\alpha$.
In Figure 1, we represent the cases when $n$ is even or odd. The rotation $R$ sends each level to the one immediately above it, except the top levels. When $x$ is in one of the tops of the towers, $R$ sends $x$ to the point $R^{q_{n}}x$ or $R^{q_{n+1}}x$ that is in the base. This means that the first return time of $R$ to the base of the Rokhlin tower is a rotation by $(-1)^{n}\|q_{n+1}\alpha\|$ $\mod\|q_{n+1}\alpha\|+\|q_{n}\alpha\|$.
Remark 9.
Notice that by adding $z\mod 1$ to the basis of towers, we can build Rokhlin towers around $z$, i.e. the basis now are $[-\|q_{n}\alpha\|+z,z)$ and $[z,\|q_{n+1}\alpha\|+z)$ or $[-\|q_{n+1}\alpha\|+z,z)$ and $[z,\|q_{n}\alpha\|+z)$ for $n$ even or odd respectively. This is convenient if we want to avoid discontinuities of the cocycle function inside the levels of the towers.
3. Skew products and ergodicity criterion
In this section, we recall the main definitions and a few results that we will use later.
3.1. Essential values of a cocycle
Let $T:X\to X$ a map and $f:X\to\mathbb{R}$ a function. We denote
$S_{n}f(x)=\left\{\begin{array}[]{lll}\sum_{k=0}^{n-1}f(T^{k}x)&\text{ if }&n\geq 1\\ 0&\text{ if }&n=0\\ -\sum_{k=1}^{|n|}f(T^{-k}x)&\text{ if }&n\leq-1\text{ and $T$ is invertible}\end{array}\right.$
Let $\mathcal{B}$ be a $\sigma$-algebra on $X$. Suppose that $T$ is ergodic, invertible, measure preserving with respect to the Borel measure $\mu$ on $\mathcal{B}$; we assume that $\mu$ is finite. Consider $f:X\to\mathbb{R}$ a measurable function and the skew product $T_{f}:(x,t)\mapsto(Tx,t+f(x))$. We say that $a\in\mathbb{R}$ is an essential value of $f$ if for all $B\in\mathcal{B}$ with $\mu(B)>0$ and for all $\epsilon>0$, there exists $n\in\mathbb{N}$ such that
$\mu\left(\{x:T^{-n}B\cap B\cap\{x:|S_{n}f(x)-a|<\epsilon\}\}\right)>0.$
Denote $Ess(T_{f})$ the set of essential values of $f$. The following is true about the previous setup:
Proposition 10 ([Aar97, 8.2.1 Proposition and 8.2.5 Corollary]).
Ghost text
  1. (1)
    $Ess(T_{f})$ is a closed subgroup of $\mathbb{R}$.
  2. (2)
    $T_{f}$ is ergodic with respect to $\mu\otimes{\rm Leb}_{\mathbb{R}}$ if and only if $Ess(T_{f})=\mathbb{R}$.
We will use the following version of an essential value criterion, as introduced by Conze and Frączek.
Theorem 11 ([CF11, Lemma 2.7]).
Let $a\in\mathbb{R}$. Assume that for every $\epsilon>0$ there exists a sequence of subsets $\{\Xi_{n}\}_{n\in\mathbb{N}}$ and an increasing sequence $\{q_{n}\}_{n\in\mathbb{N}}$ of natural numbers such that
  1. (1)
    $\liminf_{n\to\infty}\mu(\Xi_{n})>0$,
  2. (2)
    $\lim_{n\to\infty}\sup_{x\in\Xi_{n}}|T^{q_{n}}(x)-x|=0$,
  3. (3)
    $\lim_{n\to\infty}\mu(\Xi_{n}\triangle T(\Xi_{n}))=0$,
  4. (4)
    $\sup_{x\in\Xi_{n}}|S_{q_{n}}f(x)-a|<\epsilon$.
Then $a\in Ess(T_{f})$.
The result [CF11, Lemma 2.7] states that the topological support of the limit distribution $P:=\lim_{n\to\infty}\frac{1}{\mu(\Xi_{n})}(S_{q_{n}}f(x)|_{\Xi_{n}})_{*}\mu|_{\Xi_{n}}$ is contained in $Ess(T_{f})$ – this limit exists up to a subsequence due to tightness guaranteed by Condition (4). However, again by (4), the topological support of $P$ is contained in $[a-\epsilon,a+\epsilon]\subset\mathbb{R}$. By passing with $\epsilon$ to 0 and by the fact that $Ess(T_{f})$ is a closed subset of $\mathbb{R}$, we get $a\in Ess(T_{f})$.
Note that to prove ergodicity of a skew product on the space $[0,1)\times\mathbb{R}$, it is enough to restrict the computation to the induced map on $J\times\mathbb{R}$, where $J\subset[0,1)$ is of positive measure. It follows from the following more general facts.
Proposition 12 ([Aar97, 1.5.2 Proposition]).
Suppose that $S:X\to X$ is conservative and $A\subset X$ with $m(A)>0.$ Define $m|_{A}(C)=m(A\cap C)$
  1. (1)
    If $T$ is ergodic w.r.t $m$, then $T_{A}$ is ergodic w.r.t $m|_{A}$.
  2. (2)
    If $T_{A}$ is ergodic w.r.t $m|_{A}$ and $X=\bigcup_{n=1}^{\infty}T^{-n}A$, then $T$ is ergodic w.r.t $m$.
As an implication of Proposition 12, and the fact that every skew product with an ergodic base and 00-mean cocycle is conservative (see [Sch77, Theorem 5.5], [Aar97, 8.1.5 Corollary]), we have the following:
Corollary 13.
Let $T:=T_{\lambda,\pi}:I\to I$ be a 3-IET which is ergodic with respect to $\mu:={\rm Leb}_{I}$ and let $f:I\to\mathbb{R}$ be a measurable function with $\int f\textrm{d}\mu=0$. Denote the first return map of $T_{f}$ to the cylinder $J\times\mathbb{R}$ by $R_{\widetilde{f}}:J\times\mathbb{R}\to J\times\mathbb{R}$. If $R_{\widetilde{f}}$ is ergodic with respect to ${\rm Leb}_{J}\otimes{\rm Leb}_{\mathbb{R}}$ then $T_{f}$ is ergodic with respect to $\mu\otimes{\rm Leb}_{\mathbb{R}}$.
4. Proof of Theorem 1
4.1. Reducing $T_{f}$ to $R_{\widetilde{f}}$
To make the statement of Theorem 1 more precise, start with a minimal 3-IET $T$ and $f\in C^{1}([0,1))$ with
$f(0)=\lim_{x\to 1}f(x)=0.$ (4)
As consider in Cases 1 and 2 (i.e. Equations 2 and 3,) the first return time of the skew product $T_{f}$ to $J\times\mathbb{R}$ is of the form $R_{\widetilde{f}}$ where $R$ is a rotation on $J$ and $\widetilde{f}$ is a function with at most one discontinuity in the interior of $J$. In Case 1, the discontinuity is at $\lambda_{A}-\lambda_{C}$, whereas in Case 2, there are no discontinuities in the interior of $J$. All thanks to Equation 4. We note that any 3-IET is measurably isomorphic to its inverse. The isomorphism between $T$ and $T^{-1}$ is of the form $\mathcal{I}(x)=|I|-x$. Hence $T_{f}$ is measurably isomorphic to ${(T^{-1})}_{-f\circ\mathcal{I}}$. The isomorphism is $(x,t)\mapsto(\mathcal{I}(x),-t)$. Consequently, proving or disproving ergodicity of $T_{f}$ is enough to consider Case 2.
If $f$ has bounded variation, the induced cocycle $\widetilde{f}$ in Equation 3 also has bounded variation. Similarly, $\widetilde{f}$ is absolutely continuous if $f$ is absolutely continuous and satisfies Equation 4, see Remark 6.
[…]
  1. (1)
    $\widetilde{f}$ is an absolutely continuous function,
  2. (2)
    the derivative of $\widetilde{f}$ has bounded variation,
  3. (3)
    $\int\widetilde{f}=0$ and $\int\widetilde{f}^{\prime}=0$.
Because of the above reduction and Remark 7, the proof of Theorem 1 boils down to showing the following fact.
Proposition 14.
Let $R$ be a rotation by $\alpha$, an irrational number of Roth type, and let $\widetilde{f}$ be an absolutely continuous function on $[0,1)$ with $\int\widetilde{f}=0$, whose first derivative has bounded variation. Then $\widetilde{f}$ is a cobundary for are and in particular $R_{\widetilde{f}}$ is not ergodic.
4.2. The solution of the cohomological equation
The proof of Proposition 14 follows from:
Theorem 15 ([MMY05, Theorem A]).
Let $T$ be an IET with the Keane property and of Roth type. Let $\Phi\in BV^{1}_{*}\left(\bigsqcup_{s\in\mathcal{A}}I_{s}\right)$. There exists a function $\chi$ constant on each interval $I_{s}$ and a bounded variation function $\Psi$ such that
$\Psi\circ T-\Psi=\Phi-\chi.$ (5)
For context, $\Phi\in BV^{1}_{*}(\bigsqcup I_{s})$ if $\Phi|_{I_{s}}$ is absolutely continuous, $\Phi^{\prime}|_{I_{s}}$ has bounded variation and $\int\Phi^{\prime}=0.$
Proof of Proposition 14.
In the hypothesis of Theorem 15, let $T=R$ a rotation by the irrational number $\alpha$ of Roth type and let $\Phi=\widetilde{f}$. Then, there exists a piecewise constant function on $[0,1-\alpha)$ and $[1-\alpha,1)$ and a bounded variation function $\Psi$ such that
$\Psi\circ R-\Psi=\widetilde{f}-\chi.$
Since $\int\widetilde{f}=0$, the previous equation implies that $\int\chi=0$. Because $\chi=a_{1}\mathds{1}_{[0,1-\alpha)}+a_{2}\mathds{1}_{[1-\alpha,1)}$, we can assume that $a_{1}=\alpha$ and $a_{2}=-(1-\alpha)$. Notice that the function $\chi=\alpha\mathds{1}_{[0,1-\alpha)}+(\alpha-1)\mathds{1}_{[1-\alpha,1)}$ is a coboundary (i.e. $\chi=h\circ R-h$ for $h(x):=x$):
$h(x+\alpha\mod 1)-h(x)=\left\{\begin{array}[]{lc}x+\alpha-x&x\in[0,1-\alpha),\\ x-(1-\alpha)-x&x\in[1-\alpha,1).\end{array}\right.$
In conclusion $\widetilde{f}$ is a coboundary:
$\widetilde{f}=(\Psi+h)\circ R-(\Psi+h)$
and $R_{\widetilde{f}}$ is not ergodic because sets of the form $\{(x,s):s\leq\Psi(x)+h(x)\}$ are invariant. ∎
4.3. Ergodic examples and the concluding argument
[…]
$a_{j}=\lim_{x\to\beta_{j}^{+}}g(x)-\lim_{y\to\beta_{j}^{-}}g(y)$
and denote by $S(g)$ the sum of the jumps:
$S(g)=\sum_{j=0}^{k}a_{j}=-\int f^{\prime}.$
We set some notation first. We say that $g\in C_{0}^{1+\textrm{BV}}$ if $g:\mathbb{T}\to\mathbb{R}$ is absolutely continuous with $\int g=0$ and $g^{\prime}$ has bounded variation. We say that $g\in C_{0}^{1+\textrm{PAC}}$ if $g\in C_{0}^{1+\textrm{BV}}$ and $g^{\prime}$ is piecewise absolutely continuous.
In this paper, we need the following:
Theorem 16 ([Frą00, Theorem 1.1]).
Let $R$ be a rotation by $\alpha\in S_{1}^{0}$ and let $g\in C_{0}^{1+\textrm{PAC}}$ be such that $S(g^{\prime})=0$. Let $\beta_{0},\dots,\beta_{k}\in\mathbb{T}$ be the discontinuities of $g.$ If there exists a subsequence $q_{n_{1}}<q_{n_{2}}<q_{n_{3}}<\dots$ of denominators of the continued fraction of $\alpha$ such that
$\lim_{j\to\infty}q_{n_{j}}^{2}\|q_{n_{j}}\alpha\|=0\quad\text{ and }\quad\lim_{j\to\infty}\{q_{n_{j}}\beta_{i}\}=\gamma_{i},$ (6)
where $\gamma_{i}\neq\gamma_{l}$ for $i\neq l$, $i,l=0,\dots,k,$ then $R_{g}$ is ergodic.
Frączek in [Frą00] proved the previous theorem in greater generality. We only state a simpler version for our purposes.
According to Remark 6, if $f:[0,|\lambda|)\to\mathbb{R}$ is absolutely continuous and $f(0)=\lim_{x\to|\lambda|}f(x)=0$, when Case 2 applies, we obtain from Equation 3 that $\widetilde{f}:J\to\mathbb{R}$ is absolutely continuous and $\widetilde{f}(0)=\lim_{x\to|J|}\widetilde{f}(x)$. Moreover $S(f)=S(\widetilde{f})=0$ and $S(f^{\prime})=S(\widetilde{f}^{\prime})=0$. We are ready for the proof of Theorem 1:
Proof of Theorem 1.
The full measure set 3-IETs for which $f$ is a coboundary in Theorem 1 follows directly from Proposition 14 and Remark 7.
When the Condition $S(f^{\prime})=0$ holds, Theorem 16 implies that $R_{\widetilde{f}}$ is ergodic for a dense set of 2-IETs –the Condition in Equation 6 is a dense condition. The condition 2 in Proposition 12 implies that a dense set of 3-IETs $T_{f}$ is ergodic. ∎
5. Proof of Theorem 2
In this section, we prove Theorem 2. Using the essential value criterion, we will prove that skew products are ergodic (Theorem 11). The primary tool will be the discontinuity that appears in the interior (not in the orbit of 0) of one of the intervals exchanged by $T|_{J}$, which, in our case, is a rotation.
5.1. Initial reductions.
First, let us note that we only deal with the case when $\lim_{x\to 1^{-}}f(x)\neq 0$, the proof in the other case is symmetric, up to reflection. To further simplify the computations, by multiplying by a scalar if necessary, we can assume that $\lim_{x\to 1^{-}}f(x)=1$.
Without loss of generality, we assume that Case 1 (i.e., Equation 2) holds. Then, $\widetilde{f}$, seen as a function on the circle, has three discontinuities: at 0, $\lambda_{A}-\lambda_{C}$ and $\lambda_{A}$. We reverse the problem, by considering almost every $\alpha\in[0,1)$ and almost every $\beta\in[0,1)$ and a cocycle $f$, which is a piecewise $C^{1}$-function with 00-mean, which is $C^{1}$ on $[0,1-\alpha)$, $[1-\alpha,\beta)$ and $[\beta,1)$ that has a non-trivial jump in $\beta$ (i.e. $\lim_{x\to\beta^{+}}f(x)-\lim_{x\to\beta^{-}}f(x)\neq 0$). As above, we may assume that this jump equals $-1$. We devote the rest of this section to proving that under these considerations, the skew product $R_{f}$ is ergodic, with $R$ being the rotation by $\alpha$.
Figure 2. Behavior of the Birkhoff sums $S_{k}f$ for different values of $k$. The cocycle function $f$ has discontinuities at 00, $1-\alpha$, $\beta$. The argument to prove ergodicity in Theorem 2 is to estimate the value of the jumps at specific times $k$ and combine this with Theorem 11.
5.2. Bounds of the Birkhoff sums.
Let us start by recalling the standard Denjoy-Koksma inequality. Namely, if $F:[0,1)\to\mathbb{R}$ is of bounded variation, then
$\left\|S_{q_{n}}F-q_{n}\int F\right\|_{\infty}\leq Var(F),$
where $\{q_{n}\}_{n\in\mathbb{N}}$ is a sequence of partial denominators of $\alpha$. In particular, in our case, we have
$\|S_{q_{n}}f\|_{\infty}\leq Var(f)$ (7)
and
$\|S_{q_{n}}f^{\prime}\|_{\infty}\leq Var(f^{\prime}).$ (8)
From now on, we assume that $\alpha$ is of unbounded type, meaning that if $\{a_{n}\}_{n\in\mathbb{N}}$ the is sequence of partial quotients, then there exists a subsequence $\{n_{k}\}_{k\in\mathbb{N}}$ such that
$\lim_{k\to\infty}a_{n_{k}+1}=\infty.$ (9)
We begin by showing that with the above assumption, the integrals of $S_{q_{n}}(f)$ converge quickly to 0 on properly chosen intervals.
Lemma 17.
With the assumptions above, there exists a sequence $c_{k}$ of positive numbers such that
$\sup\left\{\left|\int_{J}S_{q_{n_{k}+1}}(f)\right|:J\subset[0,1)\text{ with }\ |J|=\|q_{n_{k}}\alpha\|\right\}\leq c_{k}=O(\|q_{n_{k}+1}\alpha\|).$
Proof.
Recall that for every $n\in\mathbb{N}$, the first return map to any interval $L=[c,d)\subset[0,1)$ of length $\|q_{n+1}\alpha\|+\|q_{n}\alpha\|$ is a 2-IET (a rotation) and the discontinuity point is either $c+\|q_{n+1}\alpha\|$ or $d-\|q_{n+1}\alpha\|$. For further details, see Section 2.3. In both cases, the entire interval breaks into two Rokhlin towers of intervals (seen on the circle), where the one with the shorter base has height $q_{n}$ and the one with the longer base has height $q_{n+1}$. Without loss of generality, we assume that the former holds for all $n\in\mathbb{N}$.
Fix $k\in\mathbb{N}$ and let $J=[a,b)$ as in the assumption of the lemma. Consider an interval $J^{\prime}:=[a-\|q_{n_{k}+1}\alpha\|,b)$. Then the full circle can be decomposed into two towers, one over the interval $\hat{J}=[a-\|q_{n_{k}+1}\alpha\|,a)$ of height $q_{n_{k}}$ and the other over the interval $J$ of height $q_{n_{k}+1}$. Then, the function $g:J^{\prime}\to\mathbb{R}$ given by
$g(x):=\begin{cases}S_{q_{n_{k}}}(f)(x)&\text{if}\ x\in\hat{J};\\ S_{q_{n_{k}+1}}(f)(x)&\text{if}\ x\in J.\end{cases}$
Since $\int_{0}^{1}f=0$, we also have $\int_{J^{\prime}}g=0$. Moreover, we have
$\left|\int_{\hat{J}}S_{q_{n_{k}}}(f)\right|\leq\|q_{n_{k}+1}\alpha\|\cdot Var(f).$
Thus
$\left|\int_{J}S_{q_{n_{k}+1}}(f)\right|\leq\|q_{n_{k}+1}\alpha\|\cdot Var(f).$
Then it suffices to take $c_{k}:=\|q_{n_{k}+1}\alpha\|\cdot Var(f)$. ∎
5.3. $(\delta,\infty)$-balanced points
For every $k\in\mathbb{N}$, consider the Rokhlin towers with basis on the intervals $I^{1}_{k}$ and $I^{2}_{k}$ given by
$I^{1}_{k}=[0,\|q_{n_{k}+1}\alpha\|)\ \text{and}\ I^{2}_{k}=[\|q_{n_{k}+1}\alpha\|,\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|)$ (10)
or
$I^{1}_{k}=[0,\|q_{n_{k}}\alpha\|)\ \text{and}\ I^{2}_{k}=[\|q_{n_{k}}\alpha\|,\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|),$
depending on the parity of $n_{k}$ (see Remark 9). Without loss of generality, by passing to a subsequence if necessary, we may assume that only one of the above definitions gives bases of towers in the Rokhlin tower decomposition. Suppose it is the case specified by the intervals in (10). The reasoning for the other case is analogous.
Let $\delta\in(0,1)$. We say that $z\in[0,1)$ is $(\delta,\infty)$-balanced, if there exists a subsequence $(n_{k_{\ell}})_{\ell\in\mathbb{N}}$, a real sequence of positive numbers $(\epsilon_{\ell})_{n\in\mathbb{N}}$, which converges to 00 such that if $I^{2}_{k_{\ell}}=[a_{k_{\ell}},b_{k_{\ell}}):=[\|q_{n_{k}+1}\alpha\|,\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|)$, then
$z\in\bigcup_{j=0}^{q_{n_{k_{\ell}}+1}-1}R^{j}((\delta-\epsilon_{\ell})b_{k_{\ell}},(\delta+\epsilon_{\ell})b_{k_{\ell}})$
for every $\ell\in\mathbb{N}$. It turns out that a.e. point of $[0,1)$ is $(\delta,\infty)$-balanced.
Lemma 18.
For every $\delta\in(0,1)$, a.e. $z\in[0,1)$ is $(\delta,\infty)$-balanced.
This result is the direct corollary of the following lemma, whose proof we omit. The proof is identical to the proof of Lemma 4 in [FL06].
Lemma 19.
Fix $\delta\in(0,1)$. For every $k\in\mathbb{N}$ let $\epsilon_{k}:=(\sqrt{k})^{-1}$ and consider the towers
$C_{k}:=\bigcup_{j=0}^{q_{n_{k}+1}-1}R^{j}((\delta-\epsilon_{k})b_{k},(\delta+\epsilon_{k})b_{k}).$
Then
$\limsup_{k\to\infty}C_{k}=[0,1).$
5.4. Main estimates and concluding arguments.
We now go back to the proof of Theorem 2. Let $\beta\in[0,1)$ be such that it satisfies the conclusion of Lemma 18 with $\delta=\tfrac{1}{\sqrt{2}}$, i.e. $\beta$ is $\left(\tfrac{1}{\sqrt{2}},\infty\right)$-balanced. To avoid excess notation, we write that $\beta$ is $\left(\tfrac{1}{\sqrt{2}},\infty\right)$-balanced with the sequences $(n_{k})_{k\in\mathbb{N}}$ and $(\epsilon_{k})_{k\in\mathbb{N}}$. Moreover, let $(c_{k})_{k\in\mathbb{N}}$ be the sequence given by Lemma 17. Then for every $k\in\mathbb{N}$ and every $0\leq j<q_{n_{k}+1}$, we have
$\left|\int_{R^{j}[\|q_{n_{k}+1}\alpha\|,\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|)}S_{q_{n_{k}+1}}(f)\right|\leq c_{k}=O(\|q_{n_{k}+1}\alpha\|).$ (11)
To finish the proof, we will find two sequences of subsets $(\Xi_{k}^{1})_{k\in\mathbb{N}}$ and $(\Xi_{k}^{2})_{k\in\mathbb{N}}$ of $[0,1)$ which satisfy Theorem 11 for $a$ equal to $1-\tfrac{1}{\sqrt{2}}$ and $-\tfrac{1}{\sqrt{2}}$, respectively.
For any $k\in\mathbb{N}$, define $0\leq m_{k}<q_{n_{k}+1}$ to be such that
$\beta\in R^{m_{k}}[\|q_{n_{k}+1}\alpha\|,\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|)$
and let $d_{j}^{k}=R^{j}(R^{-{m_{k}}}\beta)$ for $0\leq j\leq m_{k}$ and $d_{j}^{k}=R^{j}(R^{-{m_{k}}}\beta+\|q_{n_{k}+1}\alpha\|)$ for $m_{k}<j<q_{n_{k}+1}$. Then for every $0\leq j<q_{n_{k}+1}$ we define
$J_{j}^{1}:=[R^{j}\|q_{n_{k}+1}\alpha\|,d_{j}^{k})\ \text{and}\ J_{j}^{2}:=[d_{j}^{k},R^{j}(\|q_{n_{k}}\alpha\|+\|q_{n_{k}+1}\alpha\|)).$
Note that $S_{q_{n_{k}+1}}(f)$ is continuous on every interval of the form $J_{j}^{1}$ and $J_{j}^{2}$. Now we take
$\Xi_{k}^{1}:=\bigcup_{j=0}^{q_{n_{k}+1}-1}J_{j}^{1}\ \text{and}\ \Xi_{k}^{2}:=\bigcup_{j=0}^{q_{n_{k}+1}-1}J_{j}^{2}.$ (12)
By the choice of $\delta$, we have
$\min\left\{\liminf_{k\to\infty}{\rm Leb}(\Xi_{k}^{1}),\ \liminf_{k\to\infty}{\rm Leb}(\Xi_{k}^{2})\right\}>0.$
Since $(q_{n_{k}+1})_{k\in\mathbb{N}}$ is a subsequence of partial denominators we have the limits
$\lim_{n\to\infty}\sup|R^{q_{n_{k}+1}}(x)-x|=\lim_{n\to\infty}\|q_{n_{k}+1}\alpha\|=0.$
Moreover, both $\Xi_{k}^{1}$ and $\Xi_{k}^{2}$ are unions of at most two towers with widths smaller than $\|q_{n_{k}}\alpha\|$, we have
$\lim_{k\to\infty}{\rm Leb}(\Xi_{k}^{1}\triangle T\Xi_{k}^{1})=\lim_{k\to\infty}{\rm Leb}(\Xi_{k}^{2}\triangle T\Xi_{k}^{2})=0.$
Hence, it remains to check whether the condition 4 from Theorem 11 is satisfied with $a=1-\tfrac{1}{\sqrt{2}}$ for $(\Xi_{k}^{1})_{k\in\mathbb{N}}$ and with $a=-\tfrac{1}{\sqrt{2}}$ for $(\Xi_{k}^{2})_{k\in\mathbb{N}}$.
Fix $\epsilon>0$. Take $k\in\mathbb{N}$ and $10\leq j\leq q_{n_{k}+1}-$ 1. In view of Equation 11, we have
$\Big{|}\int_{J_{j}^{1}}S_{q_{n_{k}+1}}(f)\ +\ \int_{J_{j}^{2}}S_{q_{n_{k}+1}}(f)\Big{|}<c_{k}.$
Let $\omega_{j}^{-}:=\lim_{x\to{d_{j}^{k}}^{-}}S_{q_{n_{k}+1}}(f)(x)$ and $\omega_{j}^{+}:=\lim_{x\to{d_{j}^{k}}^{+}}S_{q_{n_{k}+1}}(f)(x)$. Since $\beta$ is chosen to be the only discontinuity of $f$ in the interior of the tower $\bigcup_{i=0}^{q_{n_{k}+1}-1}R^{j+i}[\|q_{n_{k}+1}\alpha\|,\|q_{n_{k}+1}\alpha\|+\|q_{n_{k}}\alpha\|)$ (the other discontinuities are 00 and $R^{-1}(0)$) and the jump in $\beta$ is $-1$, we have
$\omega_{j}^{+}-\omega_{j}^{-}=-1\ \text{for every }0\leq j<q_{n_{k}+1}$ (13)
Since for every $0\leq j<q_{n_{k}+1}$, the function $S_{q_{n_{k}+1}}(f)$ is continuous on both $J_{j}^{1}$ and $J_{j}^{2}$, by mean value theorem we get for every $x\in J_{j}^{1}$ that
$|S_{q_{n_{k}+1}}(f)(x)-\omega_{j}^{-}|\leq\|q_{n_{k}}\alpha\|\cdot\sup_{y\in J_{j}^{1}}|S_{q_{n_{k}+1}}(f^{\prime})(y)|\leq\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime}),$ (14)
and for every $x\in J_{j}^{2}$
$|S_{q_{n_{k}+1}}(f)(x)-\omega_{j}^{+}|\leq\|q_{n_{k}}\alpha\|\cdot\sup_{y\in J_{j}^{2}}|S_{q_{n_{k}+1}}(f^{\prime})(y)|\leq\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime}).$ (15)
Thus we have
$|J_{j}^{1}|(\omega_{j}^{-}-\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime}))\leq\int_{J_{j}^{1}}S_{q_{n_{k}+1}}(f)\leq|J_{j}^{1}|(\omega_{j}^{-}+\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime}))$
and
$|J_{j}^{2}|(\omega_{j}^{+}-\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime}))\leq\int_{J_{j}^{2}}S_{q_{n_{k}+1}}(f)\leq|J_{j}^{2}|(\omega_{j}^{+}+\|q_{n_{k}}\alpha\|\cdot Var(f^{\prime})).$
By combining the two above inequalities, we get
$|J_{j}^{1}|\omega_{j}^{-}+|J_{j}^{2}|\omega_{j}^{+}\leq c_{k}+2\|q_{n_{k}}\alpha\|^{2}\cdot Var(f^{\prime})$ (16)
and
$|J_{j}^{1}|\omega_{j}^{-}+|J_{j}^{2}|\omega_{j}^{+}\geq c_{k}-2\|q_{n_{k}}\alpha\|^{2}\cdot Var(f^{\prime}).$ (17)
Take $\varepsilon>0$. Since $c_{k}$ is of order $O(\|q_{n_{k}+1}\alpha\|)$ and $\lim_{k\to\infty}a_{n_{k}+1}=\infty$, $c_{k}$ is of order $o(\|q_{n_{k}}\alpha\|)$. Moreover, since $\beta$ is $(\tfrac{1}{\sqrt{2}},\infty)$-balanced, by taking larger $k$ if necessary, for every $0\leq j<q_{n_{k}+1}$ we have $|J_{j}^{1}|\geq\frac{1}{2}\|q_{n_{k}}\alpha\|$. By taking again larger $k$ if necessary, Equations 16 and 17 together yield
$\left|\,\omega_{j}^{-}+\frac{|J_{j}^{2}|}{|J_{j}^{1}|}\omega_{j}^{+}\right|\leq\varepsilon.$
Moreover, since $\omega_{j}^{+}-\omega_{j}^{-}=-1$, we get
$\left|\left(1+\frac{|J_{j}^{2}|}{|J_{j}^{1}|}\right)\omega_{j}^{-}-\frac{|J_{j}^{2}|}{|J_{j}^{1}|}\right|\leq\varepsilon.$ (18)
Since $\lim_{k\to\infty}\epsilon_{k}=0=\lim_{k\to\infty}\tfrac{\|q_{{n_{k}}+1}\alpha\|}{\|q_{{n_{k}}}\alpha\|}$, by increasing $k$ again if necessary, it follows from $\beta$ being $(\frac{1}{\sqrt{2}},\infty)$-balanced that
$\max_{0\leq j<q_{n_{k}+1}}\left|\frac{|J_{j}^{2}|}{|J_{j}^{1}|}-(\sqrt{2}-1)\right|\leq\varepsilon.$
By combining this inequality with Equation 18, we get
$\frac{\sqrt{2}-1-2\varepsilon}{\sqrt{2}+\varepsilon}\leq\omega_{j}^{-}\leq\frac{\sqrt{2}-1+2\varepsilon}{\sqrt{2}-\varepsilon}.$
By decreasing $\varepsilon$ and increasing $k$ again, if necessary, we obtain from the above inequality and from (14) that:
$S_{q_{n_{k}+1}}(f)(x)\in(1-\tfrac{1}{\sqrt{2}}-\epsilon,1-\tfrac{1}{\sqrt{2}}+\epsilon)\ \text{for every }x\in\Xi_{k}^{1}.$
Moreover, since $\omega_{j}^{+}-\omega_{j}^{-}=-1$, we get from (15) that
$S_{q_{n_{k}+1}}(f)(x)\in(-\tfrac{1}{\sqrt{2}}-\epsilon,-\tfrac{1}{\sqrt{2}}+\epsilon)\ \text{for every }x\in\Xi_{k}^{2}.$
This gives (4) in Theorem 11 for $(\Xi_{k}^{1})_{k\in\mathbb{N}}$ and $(\Xi_{k}^{2})_{k\in\mathbb{N}}$. Hence, the assumptions of Theorem 11 are satisfied for both sequences. Since $1-\frac{1}{\sqrt{2}}$ and $-\frac{1}{\sqrt{2}}$ generate a dense subgroup of $\mathbb{R}$, this together with Proposition 12 completes the proof of Theorem 2.
5.5. Non-ergodic examples and proof of Proposition 3.
We finalize this section by proving Proposition 3. Using the reduction given by Case 1 and Case 2, i.e., Equation 2 and Equation 3, we start by considering a skew product of a rotation and a function given by the previous cases. Assume that the rotation $R$ on $[0,1)$, obtained as an image of the Rauzy-Veech induction from the $3$-IET $T$, is of Roth type and that the discontinuity $\beta$ of $\widetilde{f}$ is in the point $R^{-m}(0)$ for some $m\geq 1$. Since the orbit of 00 is dense for every irrational rotation, the set of $3$-IETs we consider is dense in the space of all $3$-IETs.
Let $n\geq 0$ be such that $q_{n}\geq m$. Then, by taking $J:=[0,\|q_{n}\alpha\|+\|q_{n+1}\alpha\|)$, we have that the induced 2-IET $R|_{J}$ is of Roth type and the induced function $\hat{f}(x):=S_{r(x)}\widetilde{f}(x)$, where $r:J\to\mathbb{N}$ is the first return time to $J$, is continuous on both intervals of continuity of $R|_{J}$. Thus, because of Theorem 15, $\hat{f}$ is a coboundary for $R|_{J}$ and thus, by Remark 7, $f$ is a coboundary for $T$.